OpenStudy (anonymous):

The sum of three consecutive numbers is twelve less than six times the lowest number. What are the three numbers? A. 3, 4, 5 B. 10, 11, 12 C. 2, 3, 4 D. 5, 6, 7

OpenStudy (anonymous):

Just plug evaluate. A: 3+4+5 = 12 and 6*3-12 = 18-12=6, so this does not fit description. Continue this way.